OpenStudy (anonymous):

What is the period? \( \huge y = 10 \sin(\frac{2\pi}{365}(x-50)) \) Is the period \( \huge frac{2\pi}{\frac{2\pi}{365}} \) which would be 365??

OpenStudy (anonymous):

What is the period? \( \huge y = 10 \sin(\frac{2\pi}{365}(x-50)) \) Is the period \( \huge \frac{2\pi}{\frac{2\pi}{365}} \) which would be 365??

OpenStudy (anonymous):

@phi

OpenStudy (phi):

if you manipulate the coeff of x into the form \[ \frac{2\pi}{T} x \] then T is the period

OpenStudy (phi):

you have 2pi/365 the 365 matches T so yes, the period is 365

OpenStudy (anonymous):

Ok, so 365 is T so does the 50 compress the period?

OpenStudy (anonymous):

Or is shifts it to the left

OpenStudy (phi):

no the -50 * 2pi/365 is a phase shift (moves the curve to the right)

OpenStudy (anonymous):

Ok thanks so we have T = 365 with a phase shift to the right

OpenStudy (phi):

the red curve is the sin without the -50 term the black curve is yours.

OpenStudy (anonymous):

Thank you!